OpenStudy (anonymous):

i am having trouble with the algebraic part. click the question so you know the equation.

OpenStudy (anonymous):

\[\int\limits_{1}^{4} \pi (1/x-4)^{2} dx\]

OpenStudy (anonymous):

upper limit = 4 ? if yes, the denominator will become 0. The fraction does not exist. I would like to confirm whether it is really 4

OpenStudy (anonymous):

yes it is. i believe.

OpenStudy (anonymous):

the answer is 3 pi over 4

OpenStudy (anonymous):

integral pi (1/(x-4)^2) dx =integral pi (1/(x-4)^2) d(x-4) =pi (-(1/(x-4)) then, I think it is sth about taking limit but I am not sure. I am sorry for that. I am not able to help you on this.

OpenStudy (anonymous):

but i cant figure out how to get that